MICROWAVE SPECTRA, DIPOLE MOMENTS AND EQUILIBRIUM STRUCTURES OF (Z)-PENT-2-ENE-4-YNENITRILE AND MALEONITRILE

\underline{SEAN A. PEEBLES} and ROBERT L. KUCZKOWSKI, Department Of Chemistry, University Of Michigan, Ann Arbor, MI 48109-1055; ROBERT J. HALTER, RYAN L. FIMMEN and ROBERT J. McMAHON, Department Of Chemistry, University Of Wisconsin, Madison, WI 53706-1396; JOHN F. STANTON, Department Of Chemistry And Biochemistry, University Of Texas At Austin, Austin, TX 78712.

The rotational spectra of 10 isotopomers of (Z )-pent-2-ene-4-ynenitrile and 6 isotopomers of maleonitrile were measured using pulsed supersonic nozzle, Fourier-transform microwave spectroscopy. The ground state rotational constants in each case were adjusted for vibration-rotation interaction effects using ab initio force field calculations. The resulting equilibrium rotational constants were in very good agreement with those from high level ab initio calculations. Stark effect measurements allowed dipole moments of 3.53(3) D and 5.32(6) D, respectively to be determined for the two molecules. The equilibrium structural parameters and comparisons with calculated values will be presented.